OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=4/(s-1)^3.

OpenStudy (anonymous):

ok so this is inverse form of shifted function so we would have \[\frac{ 4 }{ s^3 }|s->s-1\] =>\[2\frac{ 2 }{ s^3 }|s->s-1\] so laplace of that is \[2*t^{2}\] and since we had s-1 we would have \[2t^{2}e^{t}\]

OpenStudy (anonymous):

But what you wrote looks like limit, also, where did you get the e^t?

OpenStudy (anonymous):

yeah it looks so messy , i cant write it good here dunno how to write the inverse laplace notation i will try explain better wait

OpenStudy (accessdenied):

\( \mathcal{L}^{-1} \) \mathcal{L}^{-1} Also I think you were using -> for arrow, \( \to \) \to

OpenStudy (anonymous):

yeah i didnt know how to write them -> is arrow

OpenStudy (anonymous):

Thanks.

OpenStudy (accessdenied):

\( \displaystyle \mathcal{L}^{-1} \left(\frac{4}{(s - 1)^3}\right) \) This s-variable function is shifted to the right by 1 unit, observing the \(s - 1\). We can use the concept stating that a \(s - n\) shift will correspond to the t-variable function's exponential \(e^{nt}\). Also we can quickly pull out the coefficent 4 because the operation is linear. So, once we pull the exponential, \( s - 1 \to s\) and we have... \( \displaystyle \mathcal{L}^{-1} \left(\frac{4}{(s - 1)^3}\right) = 4e^{t} \mathcal{L}^{-1} \left(\frac{1}{s^3}\right) \) As stated, our inverse Laplace transform of this s-variable function is of the form \( \displaystyle \mathcal{L}^{-1} \left(\frac{n!}{s^{n+1}}\right) = t^n \) For n=2. We pull in a factor of 2 and we can simply convert that inverse Laplace transform to the t-domain. \[ \begin{align} 4e^{t} \mathcal{L}^{-1} \left(\frac{1}{s^3}\right) &= 2e^{t} \mathcal{L}^{-1} \left(\frac{2}{s^3}\right) \\ &= 2e^{t} t^2 \\ \end{align} \] As we were hoping for. :)

OpenStudy (anonymous):

http://www.youtube.com/watch?v=FxoQvGONMj0 heres a yt video has laplace and inverse laplace of shifted function example should help understand better